OpenStudy (anonymous):

4x + y = -1 3x + 2y = 8

OpenStudy (anonymous):

Are these two seperate equation?

OpenStudy (anonymous):

I don't think so, I have to find the points for the graph.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

I have 4x = -1 - 4 (-1 - 4) + 2y = 8 So far but I don't know if I'm doing it right. Or if I am doing it right what to do after that.

OpenStudy (anonymous):

So what we have here is : 4x + y = -1 3x + 2y = 8 So What i think you havta do is solve by elimation So to do this we have to do the following:

OpenStudy (anonymous):

We multiply the equation by its coefiicant

OpenStudy (anonymous):

So: -3(4x+y=-1) 4(3x+2y=8)

OpenStudy (anonymous):

Can u Solve that?

OpenStudy (anonymous):

I can try.

OpenStudy (mertsj):

Here is a more sensible approach. From the first equation, y = -1-4x Replace y in the second equation with -1-4x

OpenStudy (mertsj):

\[3x+2(-1-4x)=8\]

OpenStudy (mertsj):

\[3x-2-8x=8\]

OpenStudy (mertsj):

\[-5x=10\]

OpenStudy (mertsj):

\[x=-2\]

OpenStudy (mertsj):

\[y=-1-4x\]

OpenStudy (mertsj):

\[y=-1-4(-2)=-1+8=7\]

OpenStudy (mertsj):

Solution is (-2,7)

OpenStudy (anonymous):

Could I email you the graphs I can choose from to see which one would be correct? and thank you!

OpenStudy (anonymous):

Never Mind I think I know which one is right!

OpenStudy (mertsj):

ok